# Post-hoc Study of Climate Microtargeting on Social Media Ads with LLMs: Thematic Insights and Fairness Evaluation

### v1

@@ 4. Task Definition
To predict the targeted demographics, we employ zero-shot prompting using the most recently [^2] September 12, 2024 released by OpenAI, o1-preview model [^3] https://openai.com/index/introducing-openai-o1-preview/. This is a new large language model that uses reinforcement learning (RL) and chain of thought (COT) techniques for complex reasoning, allowing it to think through a detailed internal process before responding to users.
The prompt template for the demographic prediction task with an explanation using LLMs can be found in Figure 3. Figure 2 shows the example prompts for gender prediction. Figure 5 (in App. A) shows the example prompts for age group prediction from the climate campaign dataset.
5. Experimental Setup
In this work, we use OpenAI Playground API to run o1-preview by keeping the default parameters.
@@ 5. Experimental Setup
Table 1. Accuracy and Misclassifications for Demographics.
5.1. Results
[…]
For baseline comparison, we use open sourced LLMs Llama 3 (llama3-70b-8192444https://github.com/meta-llama/llama3) (Touvron et al., 2023) and Mistral Large 2 (mistral-large-2407555https://mistral.ai/news/mistral-large-2407/) (Jiang et al., 2023). OpenAI’s o1-preview model outperforms the baselines both in gender and age group predictions (Table 2).
Model Demo. Acc. (%) Macro Avg. F1 (%)
o1-preview gender 90.57 90.35
o1-preview age 85.95 71.00
@@ 5.1. Results
Table 2. Baseline comparisons.
[…]
Early Working: High precision (0.950.950.95) and recall (0.900.900.90), indicating strong performance in identifying individuals in this age group.
Late Working: Moderate precision (0.500.500.50) and high recall (0.750.750.75), suggesting that while the model captures most of the actual instances, it has a higher rate of false positives.
Senior: Perfect precision (1.001.001.00) but low recall (0.330.330.33), meaning the model is very accurate when it predicts the Senior class but misses a significant number of actual senior instances.
Young: Good precision (0.760.760.76) and high recall (0.880.880.88), showing reliable performance in identifying younger individuals.
5.2. Error Analysis
Table 10 in the App. B.3 presents an analysis of ad misclassifications based on gender and age group predictions. Each entry includes the actual demographic, the predicted demographic, and a brief explanation generated by LLMs. Table 10 and explanations highlight how specific patterns and themes within an ad can lead to demographic misclassifications. In some cases, traditional gender roles and age-related interests played a significant role in these misclassifications. Understanding these nuances can help in refining predictive models and improving the accuracy of demographic targeting in future ad campaigns.
Figure 4. Prompt template for generating theme and aspects from predictions and explanations (shown as zero-shot). Inputs are shown in blue, and outputs are shown in red.
6. Ad Content Analysis
To determine which words or phrases are most commonly associated with certain demographics (answering RQ2), we identify the top-5 most frequent bigrams (two-word pairs) and trigrams (three-word pairs) of the ad content for male, female, young adults, and early working age group. Due to the small sample size, we do not show this analysis for the late working age group and senior citizens. Table 11 in App. C details the results.
These frequent bigrams and trigrams highlight key themes in the ads, particularly focusing on the phrases related to climate action, urgency (emergency), and leadership for young adults. On the other hand, frequent bigrams and trigrams of early working age group highlight the central themes of climate change, clean energy, and calls to action in the ads. Moreover, for male, these frequent bigrams and trigrams highlight the recurring themes related to climate change, clean energy, and specific campaigns or initiatives in the ads. In contrast, for female, these frequent bigrams and trigrams emphasize recurring themes around climate change, clean energy, and the ‘Build Back Better’ agenda in the ads (Table 11 in App. C ).
To calculate the statistical significance, we perform Chi-Square Test (Cochran, 1952) for Independence between male and female demographics and their bigrams/trigrams. The p-values for both the bigrams and trigrams are greater than the common significance level of 0.050.050.05. This suggests that there is no statistically significant association between gender (male and female) and the frequency of the bigrams and trigrams analyzed. Therefore, we fail to reject the null hypothesis and conclude that the observed differences in bigram and trigram frequencies between males and females could be due to chance. The same trend is noticed while performing a Chi-Square Test for Independence between the young adults (18-24) and early working (25-44) age groups and their bigrams/trigram.
@@ 6. Ad Content Analysis
Male Perceived Interests and Roles Interest in Technology and Innovation Focus on Economic and Financial Issues Property and Land Management Traditional Male Activities Engagement in Political and Infrastructure Topics Conservative Views and Skepticism
@@
Female Roles as Caregivers, Environmental Advocates, and Socially Conscious Individuals Parental and Caregiving Roles Environmental Consciousness Social Welfare and Community Involvement Empathy and Emotional Appeal Female Empowerment and Leadership Health and Safety Concerns
Table 3. Gender based Themes and Aspects of Explanations.
7. Thematic Insights of Explanations
@@
As LLMs provide explanations to provide reasoning behind their prediction, we use those explanations for thematic analysis to answer RQ3. In this analysis, we only include the correct predictions and their explanations.
7.1. Themes and Aspects of Gender Explanations
@@
We prompt LLMs in a zero-shot manner to provide the common theme and aspects under specific theme of the explanations from 404040 correct male predictions and 565656 correct female predictions. The prompt template is shown in Figure 4. We detail the theme of the gender explanation and aspects of the explanation in Table 3.
@@
7.2. Themes and Aspects of Age Explanations
We prompt LLMs in a zero-shot manner to provide the common theme and aspects under specific theme of the explanations from 222222 correct young adult predictions, 757575 correct early working age group predictions, 666 correct late working age group predictions, and 222 correct senior citizen predictions. We detail the theme of the age group explanation and aspects of that explanation in Table 12 in App. D.
From $1^{st}$ row of Table 12 (App. D), we observe that the overall theme revolves around the activism and the environmental consciousness of young adults, positioning them as a key demographic for campaigns and initiatives focused on climate change and sustainability. The explanations consistently highlight the following aspects:
@@ 7.2. Themes and Aspects of Age Explanations
From $2^{nd}$ row of Table 12 in App. D, we can see that the overall theme revolves around the proactive and responsible mindset of early working-age adults, who are not only financially capable but also motivated by a strong sense of social and environmental responsibility. They are seen as key targets for initiatives that combine sustainability with practical, career-oriented, and family-focused benefits. The explanations consistently emphasize the following aspects:
@@
From $3^{rd}$ row of Table 12 in App. D, we can notice that the overall theme revolves around the responsibilities and concerns of individuals in the late working (45-64) age group, focusing on their roles as homeowners, voters, and economically engaged citizens who are likely to be influenced by environmental, economic, and policy-related messaging. The explanations specifically emphasize the following aspects:
@@
From $4^{th}$ row of Table 12 (App. D), we notice that the overall theme centers on health and safety concerns that are particularly important to senior citizens, with a focus on programs that cater to their specific needs and the heightened risks they face in certain situations. The key aspects highlighted in the explanations are:
@@
8. Fairness and Bias Analysis
In this section, we present a comprehensive fairness analysis of the model for gender and age group classifications to answer RQ4. We evaluate the models using established fairness metrics such as Demographic Parity, Equal Opportunity, and Predictive Equality to assess their performance across different groups. By analyzing confusion matrices and classification reports, we identify any disparities in prediction accuracy and error rates between groups. Our analysis aims to identify biases, investigate the underlying reasons for any observed biases. The insights gained from this analysis are critical for guiding future research in developing fair and inclusive algorithms.
Predicted Gender Demographic Parity Ratio
Female 1.0678
Male 0.9149
Table 4. Demographic Parity for Gender.
Predicted Age Group Demographic Parity Ratio
Early Working 0.95
Late Working 1.50
Senior 0.33
Young 1.16
Table 5. Demographic Parity for Age Group.
Age Group Equal Opportunity (TPR) Predictive Equality (FPR)
Early Working 0.90 0.10
Late Working 0.75 0.05
Senior 0.33 0.00
Young 0.88 0.07
Table 6. Equal Opportunity (True Positive Rate) and Predictive Equality (False Positive Rate) for Age Groups.
Ad Text Prediction Explanations by LLMs
Table 7. Misclassified Senior Instances.
8.1. Fairness Analysis on Gender Prediction
Demographic parity examines whether each gender group receives positive predictions at equal rates. Table 4 presents the Demographic Parity ratios for each gender. A ratio of 111 indicates perfect parity. The results show that females have a slightly higher likelihood of receiving positive predictions compared to males, suggesting a minor imbalance favoring the female class.
Equal Opportunity (True Positive Rate) focuses on the True Positive Rates (TPR) across gender groups, measuring the model’s ability to correctly identify positive instances within each group. We have Female TPR: $\mathbf{0.95}$ and Male TPR: $\mathbf{0.85}$. The TPR for females is higher by 0.100.100.10, indicating that the model is more effective at correctly identifying females than males.
8.2. Fairness Analysis on Age Group Prediction
Results in Table 5 show that the late working group has a higher likelihood of receiving positive predictions, while the senior group has a significantly lower rate, indicating potential bias against seniors. The senior group has a Demographic Parity ratio of 0.330.330.33, significantly lower than the ideal value of 111, suggesting under representation in positive predictions.
Table 6 shows that the TPR for the senior group is notably lower, suggesting the model is less effective at correctly identifying individuals in this age group. The senior group has an FPR of 0, indicating no false positives, while the early working group has the highest FPR among the groups.
In Table 7, the fourth instance shows a misclassification where the model predicts it as a late working (45-64 years) age group. The ad mentions ‘working families’, which is a term commonly associated with individuals in the Late Working age group who are actively engaged in the workforce and supporting families. The content revolves around a political campaign emphasizing the need for change and active participation, themes often associated with the 456445-6445 - 64 age demographic who are typically more politically active and influential.
9. Conclusion
Our work advances the discourse on climate microtargeting by demonstrating the utility of LLMs in accurately detecting and explaining targeted messaging strategies on social media. We provide a broader theme and various aspects under each theme based on the explanations from LLMs, which we hope will be an important contribution to the CSS community. Besides, the fairness analysis conducted in our study underscores the importance of evaluating and addressing the biases. Disparities in prediction accuracy and error rates, particularly in underrepresented groups, highlight the need for more inclusive and equitable targeting methods. Ultimately, this study lays the groundwork for future investigations into the role of AI in enhancing the efficacy, transparency, and accountability of digital climate advocacy.
10. Limitations
Our analysis relies on OpenAI o1-preview model. We chose o1-preview instead of the open-source counterparts due to computational resource constraints. We show our analysis on climate campaigns dataset, but our approach can easily be adapted in any dataset. It is designed to be scalable without any modifications.
11. Ethics Statement
To the best of our knowledge, we did not violate any ethical code while conducting the research work described in this paper. We report the technical details for the reproducibility of results. In this paper, we did not introduce any new dataset; instead, we experimented using existing dataset that are adequately cited. The author’s personal views are not represented in any qualitative results we report, as it is solely an outcome derived from a machine learning model.
Appendix A Prompting
Figure 5 shows the example prompts for age group prediction from the climate campaign dataset.
Figure 5. Prompt examples for age group prediction (shown as zero-shot). (a) young adults (18-24), (b) early working age (25-44), (c) late working age (45-64). Inputs are shown in blue, and outputs are shown in red.
Appendix B Experiment Details
In this section, we provide experimental results.
B.1. Confusion Matrix
Figure 6 shows the confusion matrices for the targeted demographic group. Figure 6(a) shows the confusion matrix for gender, and Figure 6(b) shows the confusion matrix for age group prediction.
(a) gender.
(b) age
Figure 6. Confusion matrices.
B.2. Demographic Classification Report
Table 8 provides the classification report with detailed performance metrics for each gender class. Table 9 shows the classification report with detailed performance metrics for each age group.
Gender Precision Recall F1-Score Support
Female 0.89 0.95 0.92 59
Male 0.93 0.85 0.89 47
Accuracy 0.91 106
Macro Avg 0.91 0.90 0.90 106
Weighted Avg 0.91 0.91 0.91 106
Table 8. Classification Report for Gender.
Age Group Precision Recall F1-Score Support
Early Working 0.95 0.90 0.92 82
Late Working 0.50 0.75 0.60 8
Senior 1.00 0.33 0.50 6
Young 0.76 0.88 0.81 25
Accuracy 0.86 121
Macro Avg 0.80 0.72 0.71 121
Weighted Avg 0.88 0.86 0.86 121
Table 9. Classification Report for Age Group.
Ad Text Actual Prediction (Misclassification) Explanations by LLMs
Table 10. Error Analysis.
B.3. Demographic Misclassifications
Appendix C Bigrams and Trigrams
Table 11 shows the top-5 most frequent bigrams (two-word pairs) and trigrams (three-word pairs) of the ad content for male, female, young adults, and early working age groups.
Demo. Indicator Groups Top 5 Bigrams: # Occurrences Top 5 Trigrams: # Occurrences
Young adults (18-24) “bold climate” - 10 occurrences “climate emergency” - 10 occurrences “declare climate” - 10 occurrences “climate leaders” - 6 occurrences “emergency need” - 6 occurrences “declare climate emergency” - 10 occurrences “bold climate leaders” - 6 occurrences “climate emergency need” - 6 occurrences “climate leaders protect” - 6 occurrences “emergency need bold” - 6 occurrences
Age group Early working (25-44) “climate change” - 10 occurrences “fight climate” - 7 occurrences “clean energy” - 6 occurrences “america clean” - 3 occurrences “energy future” - 3 occurrences “fight climate change” - 7 occurrences “america clean energy” - 3 occurrences “clean energy future” - 3 occurrences “california treestoglobe org” - 2 occurrences “fueling america clean” - 2 occurrences
Male “climate change” - 5 occurrences “clean energy” - 4 occurrences “10 million” - 3 occurrences “million trees” - 3 occurrences “carbon emissions” - 2 occurrences “10 million trees” - 2 occurrences “carbon emissions 50” - 2 occurrences “clean energy corridor” - 2 occurrences “don live way” - 2 occurrences “torched earth ale” - 2 occurrences
Gender Female “climate change” - 7 occurrences “build back” - 6 occurrences “back better” - 6 occurrences “clean energy” - 5 occurrences “protect your” - 4 occurrences “build back better” - 6 occurrences “protect your kids” - 4 occurrences “affordable child care” - 3 occurrences “fighting climate change” - 3 occurrences “climate change and” - 3 occurrences
Table 11. Top-5 Most Frequent Bigrams and Trigrams Across Different Demographic Indicators and Groups.
Appendix D Thematic Analysis of Age Group Explanation
Table 12 details the theme of the age group explanation and aspects of that explanation.
Age group Theme of Explanation Aspects of Explanation
Young adults (18-24) Activism and Environmental Consciousness Passion for Climate Action Support for Bold Environmental Leadership Engagement with Activism Desire for Immediate Change Participation in Training and Advocacy
Early working (25-44) Proactive and Responsible Mindset Environmental Consciousness Financial Stability and Disposable Income Parental and Future Concerns Career Engagement and Professional Roles Interest in Innovation and Technology Social and Political Engagement
Late working (45-64) Responsibilities and Concerns Economic and Environmental Responsibility Homeownership and Financial Stability Voter and Policy Engagement Economic Concerns
Senior (65+) Health and Safety Concerns Health and Wellness Programs Vulnerability and Safety
Table 12. Age group based Themes and Aspects of Explanations.

### v2

@@ 4 Experimental Setup
To identify the targeted demographics, we use the most recently [^3] September 12, 2024 released by OpenAI, o1-preview model [^4] https://openai.com/index/introducing-openai-o1-preview/. This is a new large language model that uses reinforcement learning (RL) and chain of thought (COT) techniques for complex reasoning, allowing it to think through a detailed internal process before responding to users. We use OpenAI Playground API to run o1-preview by keeping the default parameters. The prompt template for the demographic prediction task with an explanation using LLMs can be found in Fig. 3 in App. A.1. Fig. 2 shows the example prompts for gender prediction. Fig. 4 in App. A.1 shows the example prompts for age group prediction from the climate campaign dataset.
@@
Table 1: Accuracy and Misclassifications for Demographics.
4.1 Results
Table 1 provides the overall accuracy of the targeted demographic prediction task by LLMs as well as a detailed breakdown of correct and incorrect predictions across each demographic category. LLMs can predict the targeted demographics with an accuracy of $88.55\%$ answering RQ1. LLMs achieve high accuracy in predicting both females ($94.92\%$) and males ($85.10\%$). A small number of females are misclassified as males, and a few males are misclassified as females. Table 1 shows high accuracy for young ($88.00\%$) and early working ($91.46\%$) categories. Performance drops for the late working ($75\%$) age group and significantly for senior ($33.33\%$) categories.
For baseline comparison, we use open sourced LLMs Llama 3 (llama3-70b-8192555https://github.com/meta-llama/llama3) Touvron et al. (2023) and Mistral Large 2 (mistral-large-2407666https://mistral.ai/news/mistral-large-2407/) Jiang et al. (2023). For llama3-70b-8192 and mistral-large-2407, we keep their default parameters. To run llama3-70b-8192, we use Groq API [^7] https://wow.groq.com/. For tokenizer, we use Meta-Llama-3-70B-Instruct from Hugging Face. For running Mistral Large 2, we use Mistral AI API. Besides, we include smaller pre-trained language model BERT Devlin et al. (2019) (bert-base-uncased) for comparison. Furthermore, we compare with simple Logistic Regression (LR) Cox (1958) trained on term frequency–inverse document frequency (tf-idf) features baseline. However, OpenAI’s o1-preview model outperforms the baselines both in gender (macro average F1 score $90.35\%$) and age group (macro average F1 score $71\%$) predictions (Table 2).
Model Demo. Acc. (%) F1 (%)
LR_tf-idf gender 69.00 65.00
LR_tf-idf age 73.00 31.00
BERT gender 72.00 70.00
BERT age 70.00 26.00
@@ 4.1 Results
o1-preview gender 90.57 90.35
o1-preview age 85.95 71.00
Table 2: Baseline comparisons. Demo. : Demographics.
@@
Male Perceived Interests and Roles Interest in Technology and Innovation, Focus on Economic and Financial Issues, Property and Land Management, Traditional Male Activities, Engagement in Political and Infrastructure Topics, Conservative Views and Skepticism
@@
Female Roles as Caregivers, Environmental Advocates, and Socially Conscious Individuals Parental and Caregiving Roles, Environmental Consciousness, Social Welfare and Community Involvement, Empathy and Emotional Appeal, Female Empowerment and Leadership, Health and Safety Concerns
Table 3: Gender based Themes and Aspects of Explanations.
4.2 Statistical Tests
We perform binomial test assessing whether the model’s gender predictions are significantly better than random guessing ($50\%$).
Null hypothesis, $H_{o}$ : probability of correct prediction is $0.5$ (random guessing)
p-value: $\sim 4.37\times 10^{-19}$
Conclusion: Reject $H_{o}$. This means the model’s accuracy is significantly better than chance, with a very strong statistical significance. Table 4 shows the statistical test results from gender subgroup.
Gender p-value Conclusion
Male $5.35\times 10^{-7}$ Reject H0
Female $5.95\times 10^{-14}$ Reject H0
Table 4: Significance tests by subgroup (gender).
Here are the results of the binomial test for age group predictions, assuming random guessing across $4$ categories ($p=0.25$):
Null hypothesis, $H_{o}$: probability of correct prediction is $0.25$ (random guessing)
p-value: $\sim 4.27\times 10^{-45}$
Conclusion: Reject $H_{o}$. The model’s performance is significantly better than random chance. This confirms that the predictions are statistically meaningful. Table 5 shows the statistical test results from age subgroup.
Age Group p-value Conclusion
Senior $4.66\times 10^{-1}$ Fail to Reject H0
Early Working $1.04\times 10^{-35}$ Reject H0
Young $5.76\times 10^{-11}$ Reject H0
Late Working $4.23\times 10^{-3}$ Reject H0
Table 5: Significance tests by subgroup (age).
4.3 Error Analysis
Table 12 in App. A.2 presents an analysis of ad misclassifications based on gender and age group predictions. Each entry includes the actual demographic, the predicted demographic, and a brief explanation generated by LLMs. Explanations highlight how specific patterns and themes within an ad can lead to demographic misclassifications. In some cases, traditional gender roles and age-related interests played a significant role in these misclassifications. Understanding these nuances can help in refining predictive models and improving the accuracy of demographic targeting in future ad campaigns.
Age group Theme of Explanation Aspects of Explanation
Young adults (18-24) Activism and Environmental Consciousness Passion for Climate Action, Support for Bold Environmental Leadership, Engagement with Activism, Desire for Immediate Change, Participation in Training and Advocacy
Early working (25-44) Proactive and Responsible Mindset Environmental Consciousness, Financial Stability and Disposable Income, Parental and Future Concerns, Career Engagement and Professional Roles, Interest in Innovation and Technology, Social and Political Engagement
Late working (45-64) Responsibilities and Concerns Economic and Environmental Responsibility, Homeownership and Financial Stability, Voter and Policy Engagement, Economic Concerns
Senior (65+) Health and Safety Concerns Health and Wellness Programs, Vulnerability and Safety
Table 6: Age group based Themes and Aspects of Explanations.
5 Thematic Insights of Explanations
@@
As LLMs provide explanations to provide reasoning behind their prediction, we use those explanations for understanding thematic insights to answer RQ2. In this analysis, we only include the correct predictions and their explanations.
5.1 Themes and Aspects of Gender Explanations
@@
We prompt LLMs to provide the common theme and aspects under specific theme of the explanations from $40$ correct male predictions and $56$ correct female predictions. The prompt template is shown in Figure 5 App. A.1. We detail the theme of gender explanation and the aspects of the explanation in Table 3. Two experts in NLP and CSS meticulously evaluate the generated themes and aspects derived from the explanations, ensuring their accuracy and alignment with the intended context. The inter-annotator agreement is $0.94$ (almost perfect agreement) using Cohen’s Kappa coefficient Cohen (1960). We detail the theme of the gender explanations and aspects of that explanation in App. A.3.
5.2 Themes and Aspects of Age Explanations
We prompt LLMs to provide the common theme and aspects under specific theme of the explanations from $22$ correct young adult predictions, $75$ correct early working age group predictions, $6$ correct late working age group predictions, and $2$ correct senior citizen predictions. We detail the theme of the age group explanation and aspects in Table 6. Similarly to gender, the same two NLP and CSS experts review the generated themes and aspects from the explanations to ensure accuracy and contextual alignment. For the age group, the inter-annotator agreement is $0.91$ (almost perfect agreement) using Cohen’s Kappa coefficient. We detail the theme of the age group explanations and aspects of that explanation in App. A.4.
5.3 Baselines
We add Latent Dirichlet Allocation (LDA) Blei et al. (2003) and BERTopic Grootendorst (2022) as unsupervised techniques for generating topics from LLMs-generated explanations. BERTopic [^8] https://maartengr.github.io/BERTopic/#common is a hierarchical topic modeling and an off-the-shelf neural baseline for clustering texts. BERTopic generates two topics from the prediction explanation of gender and age (Table 7). We run LDA with four topics and show the top $10$ words in Table 7.
Dimension Method Topics / Top Words
Gender BERTopic Topic 1: Practical and Economic Appeals to Men (Technology, Property, Costs)
Topic 2: Emotional and Social Appeals to Women (Spirituality, Motherhood, Environment)
LDA Topic 1: like, resonate, men, ads, males, focus, property, care, leave, paid
Topic 2: men, ad, male, energy, traditionally, engagement, likely, emphasizes, climate, higher
Topic 3: ad, female, environmental, emphasizes, resonate, typically, males, audiences, target, ads
Topic 4: women, ad, environmental, making, issues, future, likely, mothers, childrens, concerned
Age BERTopic Topic 1: Career-Oriented Adults Focused on Sustainability and Economic Impact
Topic 2: Young, Socially Active, and Environmentally Conscious Professionals
LDA Topic 1: climate, action, ad, likely, change, young, passionate, targets, eager, adults
Topic 2: likely, engaged, environmental, group, energy, climate, issues, adults, support, typically
Topic 3: group, likely, age, sustainability, energy, aligning, engaged, clean, career, corporate
Topic 4: group, conscious, environmentally, sustainable, making, disposable, income, likely, like, products
Table 7: Comparison of BERTopic and LDA-generated topics for demographic dimensions of Gender and Age. BERTopic provides interpretable topic labels, while LDA results include top-10 words per topic.
6 Fairness and Bias Analysis
In this section, we present a comprehensive fairness analysis of the model for gender and age group classifications to answer RQ3. We evaluate the models using established fairness metrics such as Demographic Parity, Equal Opportunity, and Predictive Equality to assess their performance across different groups. By analyzing confusion matrices and classification reports, we identify any disparities in prediction accuracy and error rates between groups. Our analysis aims to identify biases, investigate the underlying reasons for any observed biases. The insights gained from this analysis are critical for guiding future research in developing fair and inclusive algorithms.
Ad Text Prediction Explanations by LLMs
Table 8: Misclassified Senior Instances.
6.1 Fairness Analysis on Gender Prediction
Demographic parity examines whether each gender group receives positive predictions at equal rates. Table 9 presents the Demographic Parity ratios for each gender. A ratio of $1$ indicates perfect parity. The results show that females have a slightly higher likelihood of receiving positive predictions compared to males, suggesting a minor imbalance favoring the female class.
Equal Opportunity (True Positive Rate) focuses on the True Positive Rates (TPR) across gender groups, measuring the model’s ability to correctly identify positive instances within each group. We have Female TPR: $\mathbf{0.95}$ and Male TPR: $\mathbf{0.85}$. The TPR for females is higher by $0.10$, indicating that model may be more effective at correctly identifying females than males
Gender Demographic Parity Ratio
Female 1.0678
Male 0.9149
Table 9: Demographic Parity for Predicted Gender.
Age Group Demographic Parity Ratio
Early Working 0.95
Late Working 1.50
Senior 0.33
Young 1.16
Table 10: Demographic Parity for Predicted Age Group.
Age Group Equal Opportunity (TPR) Predictive Equality (FPR)
Early Working 0.90 0.10
Late Working 0.75 0.05
Senior 0.33 0.00
Young 0.88 0.07
Table 11: Equal Opportunity (True Positive Rate) and Predictive Equality (False Positive Rate) for Age Groups.
6.2 Fairness Analysis on Age Prediction
Results in Table 10 show that the late working group has a higher likelihood of receiving positive predictions, while the senior group has a significantly lower rate, indicating potential bias against seniors. The senior group has a Demographic Parity ratio of $0.33$, significantly lower than the ideal value of $1$, suggesting under representation in positive predictions.
Table 11 shows that the TPR for the senior group is notably lower, suggesting the model is less effective at correctly identifying individuals in this age group. The senior group has an FPR of 00, indicating no false positives, while the early working group has the highest FPR among the groups.
In Table 8, the fourth instance shows a misclassification where the model predicts it as a late working (45-64 years) age group. The ad mentions ‘working families’, which is a term commonly associated with individuals in the Late Working age group who are actively engaged in the workforce and supporting families. The content revolves around a political campaign emphasizing the need for change and active participation, themes often associated with the $45-64$ age demographic who are typically more politically active and influential.
6.3 Statistical Significance Tests
While a 0.10 difference in TPR may seem modest, prior fairness literature Hardt et al. (2016) recognizes it as meaningful in sensitive domains. We conduct two-proportion Z-tests to assess whether the observed differences in TPRs across demographic groups are statistically significant. For gender prediction, the difference in TPR between females ($0.95$) and males ($0.85$) resulted in a p-value of $0.086$ with a $95\%$ confidence interval of $[-0.018,0.230]$. This result indicates that the gap is not statistically significant at the conventional $0.05$ level. However, in the age group analysis, the TPR difference between young adults ($0.88$) and seniors ($0.33$) yields a p-value of $0.004$ with a $95\%$ confidence interval of $[0.138,0.796]$, indicating a statistically significant disparity. This validates our concern about fairness in age-related predictions, particularly for underrepresented senior groups.
7 Conclusion
Our work advances the discourse on climate microtargeting by demonstrating the utility of LLMs in accurately detecting and explaining targeted messaging strategies on social media. We show our analysis on climate campaigns dataset, but our approach can easily be adapted in any dataset. It is designed to be scalable without any modifications. We provide a broader theme and various aspects under each theme based on the explanations from LLMs, which we hope will be an important contribution to the CSS community. Besides, the fairness analysis conducted in our study underscores the importance of evaluating and addressing the biases. Disparities in prediction accuracy and error rates, particularly in underrepresented groups, highlight the need for more inclusive and equitable targeting methods. Ultimately, this study lays the groundwork for future investigations into the role of AI in enhancing the efficacy, transparency, and accountability of digital climate advocacy.
8 Limitations
Though we show the climate microtargeting case study, it can be applicable to political or health-related campaigns. However, the main idea of this work is to utilize LLMs as a tool to analyze real-world targeting practices, measuring biases and fairness in actual campaigns.
Our analysis relies on the OpenAI o1-preview model. We chose o1-preview instead of the open-source counterparts due to computational resource constraints. We only used pre-trained LLMs and did not consider fine-tuning due to the resource constraints.
The original dataset does not specify intersectional targeting, such as young females only. Our paper acknowledges that our approach infers targeting from exclusive impression data (view), since the Facebook Ad Library API does not disclose actual targeting parameters.
We have small sample size due to restrictive filtering criteria, might result in underrepresented groups, particularly late working and senior demographics.
9 Ethics Statement
To the best of our knowledge, we did not violate any ethical code while conducting the research work described in this paper. We report the technical details for the reproducibility of the results. In this paper, we did not introduce any new dataset; instead, we experimented using existing dataset that are adequately cited. The author’s personal views are not represented in any qualitative results we report, as it is solely an outcome derived from machine learning models.
Appendix A Appendix
A.1 Prompting
Fig. 3 illustrates the prompt template. Fig. 4 shows the example prompts for age group prediction from the climate campaign dataset. Prompt template for generating theme and aspects from predictions and explanations is shown in Fig. 5.
Figure 3: Prompt template for targeted demographic prediction. (a) gender, (b) age group. Inputs are shown in blue, and outputs are shown in red.
Figure 4: Prompt examples for age group prediction. (a) young adults, (b) early working age, (c) late working age. Inputs are shown in blue, and outputs are shown in red.
Figure 5: Prompt template for generating theme and aspects from predictions and explanations. Inputs are shown in blue, and outputs are shown in red.
A.2 Demographic Misclassifications
Ad Text Actual Prediction (Misclassification) Explanations by LLMs
Table 12: Error Analysis.
A.3 Thematic Analysis of Gender Explanations
@@
A.4 Thematic Analysis of Age Explanations
From $1^{st}$ row of Table 6, we observe that the overall theme revolves around the activism and the environmental consciousness of young adults, positioning them as a key demographic for campaigns and initiatives focused on climate change and sustainability. The explanations highlight the following aspects:
@@ A.4 Thematic Analysis of Age Explanations
From $2^{nd}$ row of Table 6, we can see that the overall theme revolves around the proactive and responsible mindset of early working-age adults, who are not only financially capable but also motivated by a strong sense of social and environmental responsibility. They are seen as key targets for initiatives that combine sustainability with practical, career-oriented, and family-focused benefits. The explanations consistently emphasize the following aspects:
@@
From $3^{rd}$ row of Table 6, we can notice that the overall theme revolves around the responsibilities and concerns of individuals in the late working (45-64) age group, focusing on their roles as homeowners, voters, and economically engaged citizens who are likely to be influenced by environmental, economic, and policy-related messaging. The explanations specifically emphasize the following aspects:
@@
From $4^{th}$ row of Table 6, we notice that the overall theme centers on health and safety concerns that are particularly important to senior citizens, with a focus on programs that cater to their specific needs and the heightened risks they face in certain situations. The key aspects highlighted in the explanations are: